a uniform disc of radius R metre and mass M kg can rotate without friction on an axle passing through it's centre and perpendicular to it's plane. a cord is round at the rim of the disc and a homogenous force of F n is applied on the cord. evaluate the tangental acceleration of a point on the rim of the disc
1
Expert's answer
2021-08-27T08:28:53-0400

The Newton's second law for rotational motion says

"I\beta=\tau"

or

"\frac{MR^2}{2}*\frac{a_{\tau}}{R}=F*R"

Hence, the tangental acceleration of a point on the rim of the disc

"a_{\tau}=2F/M"